OpenStudy (anonymous):

I really need help!! ( 10x+2)^-1

OpenStudy (anonymous):

What are you trying to do?

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

(10x+2)^-1<0

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

Find out when \[ 10x+2=0 \]first.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

Isolate \(x\).

OpenStudy (anonymous):

Subtract and then divide.

OpenStudy (anonymous):

2/-10

OpenStudy (anonymous):

Now test a number that is greater than -2/10 and test a number that is less that -2/10.

OpenStudy (anonymous):

For example, 0 is greater than it and -1 is less than it.

OpenStudy (anonymous):

Is \[ \frac 1{10x+2}<0 \]when \(x=0\)?